OpenStudy (anonymous):

Find the volume of a cylinder with a height of 2 meters and a diameter that is 5 times the measure of the height. Use 3.14 for pi and round your answer to the nearest hundredth. (5 points) Select one: a. 157 m3 b. 339.12 m3 c. 678.24 m3 d. 1570 m3

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (paki):

use formula... v=pie * r^2 * h.... and solve.,..

OpenStudy (sidsiddhartha):

Volume of a cylinder = (3.14) × r² × h

OpenStudy (anonymous):

what is the radius and the height

OpenStudy (anonymous):

@ganeshie8

OpenStudy (paki):

height here is 2 meters... and radius is 5*2= 10 = 10/2 = 5 meters... now put in the formula and solve...

OpenStudy (anonymous):

so the radius is 10 you say

OpenStudy (paki):

no... 5....

OpenStudy (anonymous):

so how would the formula look now with all of the numbers plugged in

OpenStudy (paki):

as diameter is 5 times the measure of hieght... so it will be 5*2 = 10... and radius=diameter/2.... so 10/2 = 5 meter...

OpenStudy (paki):

simple... jut put the values in the formula...and post here... so that we can help you...

OpenStudy (paki):

put your values in the formula.... v=pie * r^2 * h...

OpenStudy (anonymous):

(3.14) × r² × h (3.14) × (?)^2 × (?) what are the question marks

OpenStudy (paki):

v=pie * r^2 * h v= 3.14 * (5)^2 * 2.... height is 2 meter according to your question.... radius is 5 meter which we have calculated... and pie is 3.14.... now solve....

OpenStudy (paki):

@prowrestler now is it clear....?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 157

OpenStudy (paki):

yes... that is the answer... good :)